OpenStudy (anonymous):

A basket contains 15 pieces of fruit: 7 apples, 5 oranges, and 3 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple

OpenStudy (raden):

does Jonas return it or no ?

OpenStudy (anonymous):

5/15 * 7/14 =1/6

OpenStudy (anonymous):

Jonas's probability is 5/30 = 1/6.

OpenStudy (anonymous):

If he returns the orange, then Beth's probability is 7/30. But if he doesn't return it, then her probability is 7/29.

OpenStudy (anonymous):

it doesnt say if they return them

OpenStudy (raden):

yea, i agree with@Jas9420

OpenStudy (anonymous):

Well, then I guess we have to assume that he doesn't return the orange that he had taken.

OpenStudy (anonymous):

how did you get 1/6

OpenStudy (anonymous):

Using the Probability Formula which is P(A) = Number of ways an event can occur/ The total number of possible outcomes.

OpenStudy (anonymous):

how do you use P(A)

OpenStudy (anonymous):

The number of ways an event can occur in Jonas' case is 5 because that is the number of apples that are there. And the total number of possible outcomes would be the sum of all the fruits which is 15+7+5+3.

OpenStudy (anonymous):

P(A) is nothing but the probability.

OpenStudy (anonymous):

so the number of how much is there al ltogether and the muber of hw many of the single one is there

OpenStudy (anonymous):

Sort of.

OpenStudy (anonymous):

Suppose you have 5 blue balls, 5 red balls and 5 yellow balls. What is the possibility that you'll pick a blue ball in the first try?

OpenStudy (anonymous):

5/15 which goes dwn to 1/3

OpenStudy (anonymous):

Exactly. Now, compare this to your question.

OpenStudy (anonymous):

ok i think i got it . can u help me with one more plz

OpenStudy (anonymous):

Yeah, sure.

OpenStudy (anonymous):

A construction company orders 2900 new light bulbs for a new office building. If the probability of a light bulb being defective is 1.9%, predict how many of the light bulbs are likely to be defective. Round your answer to the nearest whole number

OpenStudy (anonymous):

Do you have the options?

OpenStudy (anonymous):

The answer should be around 16 bulbs(approx.)

OpenStudy (anonymous):

A. 55 light bulbs B. 69 light bulbs C. 46 light bulbs D. 551 light bulbs

OpenStudy (anonymous):

those are my options

OpenStudy (raden):

just multiplying 1.9% * 2900 = ...

OpenStudy (anonymous):

Ah, okay.

OpenStudy (anonymous):

It's 55.1. The answer.

OpenStudy (anonymous):

1.9 bulbs out of every 100 are defective. So how many bulbs out of 2900 are defective?

OpenStudy (anonymous):

thank you so much for the help im really bad at math

OpenStudy (anonymous):

Prestige Builders has a development of new homes. There are four different floor plans, seven exterior colors, and an option of either a two-car or a three-car garage. How many choices are there for one home?